Circular No. 8932

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 F1 (SOHO) Further to IAUC 8926, an additional apparent Meyer-group comet has

been found on SOHO website images; first appearing in images as fairly bright (mag ∼ 6) but fading fast, it was elongated in the direction of motion (but not diffuse at all) — typical for Meyer-group objects.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 F1 Mar. 19.938 0^{h}03.^{m}9 + 1^{o}22′ C2 RK 2008-F32

COMET C/2008 G1 (GIBBS) A. R. Gibbs reports his discovery of a comet (discovery observation

tabulated below) on images taken with the Mt. Lemmon 1.5-m reflector, noting that four co-added 30-s unfiltered CCD exposures show a very compact coma of diameter 4′′ and a narrow 65′′ tail in p.a. 225^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other astrometric observers have also noted the object’s cometary appearance, including L. Buzzi (Varese, Italy, 0.60-m reflector, Apr. 7.92–7.96 UT; slightly diffuse with a possible 10′′ tail in p.a. 225^{o}), L. Donato and M. Gonano (Remanzacco, Italy, 0.45-m reflector, Apr. 7.95–7.96; 100 stacked unfiltered 60-s exposures show a coma diameter of nearly 10′′ and a 20′′ tail in p.a. 230^{o}), E. Guido and G. Sostero (Castellammare di Stabia, Italy, remotely using a 0.25-m reflector near Mayhill, NM, U.S.A., Apr. 8.33–8.35; 50 co-added 60-s unfiltered exposures show a short tail nearly 15′′ long toward p.a. 190^{o} and a coma of diameter 8′′), J. G. Ries (McDonald Observatory, 0.76-m reflector, Apr. 8.41–8.42; narrow tail ≈ 20′ long toward the southwest), J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m reflector, Apr. 9.32–9.34; eight co-added 60-s exposures show a 7′′ coma and a 22′′ tail in p.a. 220^{o}), and A. C. Gilmore and P. M. Kilmartin (Mt. John 1.0-m reflector, Apr. 9.63– 9.64; asteroidal head with narrow 15′′ tail in p.a. ∼ 200^{o}).

2008 UT  α_{2000}  δ_{2000}  Mag.
Apr. 7.35552  15^{h}29^{m}45.^{s}12  +4^{o}54′32′′.1  19.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-G58.

T = 2009 Feb. 7.311 TT  ω = 71.^{o}159
Ω = 215.550 }2000.0
q = 3.72077 AU  i = 72.036
2008 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green